# AI-Driven Intrusion Detection Systems (IDS) on the ROAD dataset: A Comparative Analysis for automotive Controller Area Network (CAN)

### v1

@@ 1. Introduction
By taking into consideration that the effectiveness and robustness of an Intrusion Detection System (IDS) are not solely dependent on the capabilities of the AI model in use, but also on the characteristics of the training dataset, we initiate the process of training different models using a unique dataset (verma2022addressing) that includes 11 discrete categories of stealthy attacks, validated on a real vehicle, marking a significant departure from previous datasets, which were predominantly composed of few unrealistic attacks (lampe2023cantrainandtest). Starting by labeling (verma2022addressing), this paper’s contributions extend to the evaluation of a wide range of both supervised and unsupervised deep learning techniques, including the Transformer-based Attention Network (TAN) developed by Nguyen et al. (2023) (10141599), the Deep Convolutional Neural Network (DCNN) developed by Song et al. (2020) (SONG2020100198), and the model based on Long Short-Term Memory (LSTM) developed by Hossain et al. (2020) (Hossain2020LSTMBasedID), alongside various machine learning methods such as LightGBM, Random Forest (RF) and the LCCDE model (lccde) applied to the aforementioned dataset. In conclusion, we conduct a comparative analysis, employing additional established datasets such as the HCRL Car Hacking dataset and the In-Vehicle Network Intrusion Detection Challenge dataset, eventually suggesting improvements.
@@ 2.1.2. CAN vulnerabilities
Some approaches to secure the communication on the bus have been experimented during the years (vector-encryption), but unfortunately they have not gained enough traction. Consequently, the protocol does not respect 2 of the 3 fundamental properties focused on by information security principles, which are summarized by the CIA triad, Confidentiality, Integrity, Availability:
@@
The most difficult task for an attacker, considering the critical vulnerabilities of the CAN protocol, would be to access the bus. The main attack vectors are two (experimental-security-analysis-automobiles), one is physical access, the other one is through one of the many wireless interfaces.
@@
A challenge-response mechanism protects ECUs from being reflashed and reading sensible memory. This challenge-response is performed through a seed-to-key challenge, but both seed and key are fixed and stored in the device. The system allows only an attempt every 10 seconds, but due to the low complexity, it is sufficient to have access for 7 and a half days to be able to crack the challenge-response and therefore reflash every ECU (experimental-security-analysis-automobiles).
The CAN bus exposes also a diagnostic system called Device Control, a powerful tool that is meant to be used for testing purposes, and it can modify the internal state of the ECUs. Paired with the vulnerabilities already described, it is easily exploitable and it can lead to dangerous consequences. This service takes an argument called a Control Packet Identifier (CPID), which specifies a group of controls to override, and each one can take up to five bytes as parameters, specifying which controls in the group are being overridden, and how to override them (experimental-security-analysis-automobiles).
@@ 2.2. Literature review and research gaps
One of the first security analyses on real cars has been made in 2010 by Koscher et al. (experimental-security-analysis-automobiles), assessing the behavior of the vehicle during specific attacks. They leveraged the security vulnerabilities of CAN to control a wide range of critical functions, like stopping the engine, disabling the brakes, or even braking on-demand individual wheels. However, they do not yet prove that these attacks are feasible remotely. In 2011 Checkoway et al. (checkoway2011comprehensive) analyzed purely the external attack surface of one vehicle to assess if it might be vulnerable to remote attacks exploiting the wireless interfaces present in the car, and they found out that it is indeed possible through multiple short-range and long-range vectors like Bluetooth or cellular interfaces. The results of Checkoway et al. are particularly concerning if we take into consideration the paper published by Koscher et al.: these two studies show that intruders can preserve their anonymity controlling critical features of vehicles from long distances, potentially causing disastrous, life-threatening consequences.
@@
Charlie Miller and Chris Valasek are two researchers working on automotive cybersecurity who gained a reputation during the years by publishing multiple relevant papers on the topic. In 2013 (adventures-in-automotive-networks) they tinkered with the CAN bus, exploiting its vulnerabilities to take control of critical functionalities of two different cars; in 2014 (survey-of-remote-sufaces) they took Checkoway et al.’s previous research further by identifying the remote attack surface for a large range of vehicles, identifying the ECUs’ capabilities and estimating the difficulty of remote exploitation; in 2015 (remote-exploitation-miller-valasek) they brought to conclusion 3 years of research disclosing a long-range remote attack that can be performed on Jeep Cherokees without any physical interaction, causing Fiat Chrysler to recall 1.4 million vehicles (fca-recall). One year later, in 2016 (can-message-injection), they published another paper explaining how they performed a masquerade attack on the Jeep Cherokee and how they were able to deal with message confliction.
@@
Charlie Miller and Chris Valasek believe that attack detection represents an inexpensive and accurate way to greatly improve the security of CAN networks (survey-of-remote-sufaces). Current research on the topic focuses extensively on machine learning (bari2023intrusion) and deep learning (lin2022using) approaches in order to reach an increasingly higher accuracy. Unfortunately, these systems are not yet perfect: in 2021 Bhatia et al. (bhatia2021evading) explained how they developed a masquerade attack called DUET which could evade detection of Voltage-based Intrusion Detection Systems. This type of IDS is able to observe the voltage fingerprint of ECUs, thus recognizing the ones belonging to attackers. However, the DUET attack showed that VIDS are not a silver bullet against intrusions, so the authors introduced a lightweight defense mechanism called RAID, suggesting to modify the frame format in a protocol-compatible way in order to safeguard against the corruption of ECUs’ voltage fingerprints.
@@
According to its design, an Intrusion Detection System can be categorized into two major families: anomaly-based and signature-based detection methods (lokman2019intrusion). While signature-based IDSs compare the traffic with a database of known attack patterns, anomaly-based IDSs can be powered by machine learning techniques to model the normal behavior of the machine and they are able to detect unseen attacks as well.
@@
Although prior AI-based IDS models have yielded to promising results (lccde)(Hossain2020LSTMBasedID)(SONG2020100198)(10141599), there exists a remarkable gap in testing them on a more realistic dataset including sophisticated attack patterns. It is worth noting that the effectiveness of an IDS model depends heavily on the quality of the data it was trained on. Previously introduced approaches were tested on datasets characterized by a limited number of attack types, typically only three or four, which are considered simple and unrealistic as they are conducted in a testbed environment (lampe2023cantrainandtest). These attacks are easily detectable and not necessarily verified on a real car. To build a robust IDS able to detect stealthy attacks it is imperative to test the model against high fidelity attack scenarios.
@@ 3.1. ROAD dataset and Labeling
The Real ORNL Automotive Dynamometer (ROAD) dataset (verma2022addressing) includes 33 attack captures spanning 30 minutes and 12 ambient captures over 3 hours (Table 1). It originates from a real undisclosed vehicle driven on a dynamometer of the Oak Ridge National Laboratory (ORNL) during 3 and a half hours of activity, reproducing various benign driving behaviors, while ambient and attack data were collected.
The logs are collected with the candump utility (candump) and the output format is the following:
@@
    Targeted ID Fabrication Attacks: These attacks are performed injecting a message immediately after the legitimate message targeting a specific ID. In this case only the target signal is modified. Differently from most of public datasets, the fabrication attacks in the ROAD dataset contain one single frame injected between ambient frames of the same ID, since only a single injected message is needed after each legitimate message for this attack to be successful. Because these are real, physically verified attacks with the minimally occurring injected frames (due to the flam delivery), they provide perhaps the best (i.e., most stealthy/most difficult to detect), current, public data for testing frequency-based IDSs (verma2022addressing).
@@ 3.2. The HCRL Car-Hacking dataset
The Car-Hacking dataset (car-hacking-dataset), according to Rajapaksha et al. (2023) (rajapakshaetal), is the most popular dataset in literature, and in many recent publications it seems to be used without citation (verma2022addressing). It is generated by injecting CAN messages into the Hyundai YF Sonata’s CAN bus through the OBD-II port and includes both attack and regular messages. Each sample is composed by a timestamp, the CAN ID, the DLC and up to 8 bytes of the data field. The attack types are four:
  • DoS attack: This attack is simply performed by transmitting messages on the CAN bus with the ID containing only dominant bits (0) (i.e. ID=0x000) every 0.3 milliseconds. As seen in the previous section, the CAN bus is using wired-AND signaling, which means that the messages with the ID containing the largest number of dominant bits will be prioritized over any other ECU, which will have to patiently wait for the attack frame to be completely transmitted before trying to send another one. Therefore, an attacker with access to the CAN bus would simply need to continuously inject messages using this technique and the correct functioning of the system would be disrupted, needing a manual restart of the car (experimental-security-analysis-automobiles).
@@ 3.3. The In-Vehicle Network Intrusion Detection Challenge dataset
The In-Vehicle Network Intrusion Detection Challenge dataset (in-vehicle-dataset) was used during the ”Information Security R&D Data Challenge 2019” and includes traffic data collected from three vehicles in stationary state: a Huyndai Sonata, a Kia Soul, and a Chevrolet Spark. We have used the datasets collected during the preliminary round, since they include three attack categories (Flooding, Fuzzy, and Malfunction) compared to the two attack categories of the final round. The attack types are the same of the Car-Hacking dataset, but with different names. The dataset is labelled and, like the Car-Hacking dataset, each sample is represented by a timestamp, a CAN ID, the DLC and the data field. For all three types of vehicles, the proportion of ”Flooding” attacks is relatively high, about 20% of the total messages, compared to only about 8% for the other two types (Table 3). The application of the same attack types and setup on three different vehicles makes this dataset unique and enables us to test transfer learning approaches on it, in order to improve the generalizability of the model.
@@ 4.1. LCCDE and ML approaches
To compare neural network approaches to traditional machine learning ones, gradient boosting and decision tree based approaches have also been tested on the available datasets. Specifically we considered Random Forest, LightGBM and the LCCDE model. These methods, like other machine learning algorithms, are preferable because of their speed, efficiency and explainability. Furthermore, the Leader Class and Confidence Decision Ensemble (LCCDE) (lccde), an ensemble model based on gradient boosting algorithms, was able to reach state-of-the-art results. The ensemble algorithm is divided into two stages: training and prediction. In the training stage, the three base models are trained on the labeled data. For each class, the leader is selected based on the best F1-score of the models, obtained from cross-validation. In case of a tie in F1-score, the model with the highest speed is chosen.
@@ 4.2. LSTM-based model
LSTMs are a type of Recurrent Neural Networks (RNNs) that are frequently and effectively applied for sequence classification tasks, therefore we have selected the state-of-the-art LSTM model described by Hossain et al. (Hossain2020LSTMBasedID) to test this approach on the previously specified datasets, since it achieved outstanding results on the Survival Analysis Dataset (survival-analysis-dataset). The authors have tested many hyperparameters for both binary and multiclass classification. The best performing model was composed by a single LSTM layer with 512 units using single time steps and batches of size 512 as input. Each input sample includes ID, DLC (if available) and the bytes of the data field. We have performed multiclass classification on every dataset and binary classification only on the ROAD dataset, to compare its results with the DCNN and the TAN models. However, the approach of analyzing single messages can be considered vulnerable to replay attacks and since the authors used single time steps, the utility of a LSTM for this purpose is questionable.
Figure 2. Frame creation process of the DCNN.
@@ 4.3. DCNN
The IDS model proposed by Song et al. (SONG2020100198) is based on a convolutional neural network (CNN), aimed at detecting the patterns of IDs appearing in the CAN traffic. During the training phase, the model employs a frame builder to extract sequences of IDs from the logged traffic data. To facilitate CNN processing, which typically works with grid-formatted input, the temporal sequential patterns of CAN traffic are processed as spatially-local correlation. The frame builder extracts the 29-bit identifiers, which are the maximum number of bits for representing an ID when using the ’extended format’ (can-specification), from the most recent 29 messages and builds a $29\times 29$ bitwise frame by stacking them. A sliding window of size 29, with a step of 29 is used. This means that attack patterns might extend across 2 different frames and could not be easily recognized by the DCNN, which is unable to retain memory across frames. For this reason, we tested a different approach modifying the step of the sliding window to one, enabling the model to catch every pattern involving IDs and extending for any 29 consecutive frames. The frame creation process is illustrated in Figure 2.
@@
The DCNN model itself is based on the Inception-ResNet model (szegedy2016inceptionv4inceptionresnetimpactresidual), which is originally designed to classify $299\times 299\times 3$ input image data into 1000 classes, but has been reduced by the authors to classify more efficiently $29\times 29\times 1$ into only 2 classes. Compared to the original model the proposed DCNN excludes the Inception-ResNet-C block, while keeping the stem, Inception-ResNet-A, -B, Reduction-A and -B, average pooling, dropout, and softmax blocks, although they have been internally simplified to adapt to the studied case. The final model uses only 2% of the memory of the original one, and only the 18% of the original parameters, making it more suitable for use in restricted environments, like automotive applications.

### v2

@@ Abstract.
AIoT, Intrusion Detection System, IDS, Controller Area Network, CAN, ROAD Dataset
copyright: none
@@ 1. Introduction
By taking into consideration that the effectiveness and robustness of an Intrusion Detection System (IDS) are not solely dependent on the capabilities of the AI model in use, but also on the characteristics of the training dataset, we initiate the process of training different models using a unique dataset (Verma et al., 2024) that includes 11 discrete categories of stealthy attacks, validated on a real vehicle, marking a significant departure from previous datasets, which were predominantly composed of few unrealistic attacks (Lampe and Meng, 2023). Starting by labeling (Verma et al., 2024), this paper’s contributions extend to the evaluation of a wide range of supervised deep learning techniques, including the Transformer-based Attention Network (TAN) developed by Nguyen et al. (2023) (Nguyen et al., 2023), the Deep Convolutional Neural Network (DCNN) developed by Song et al. (2020) (Song et al., 2020), and the model based on Long Short-Term Memory (LSTM) developed by Hossain et al. (2020) (Hossain et al., 2020), alongside various machine learning methods such as LightGBM, Random Forest (RF) and the LCCDE model (Yang et al., 2022) applied to the aforementioned dataset. In conclusion, we conduct a comparative analysis, employing additional established datasets such as the HCRL Car Hacking dataset and the In-Vehicle Network Intrusion Detection Challenge dataset, eventually suggesting improvements.
@@ 2.1.1. Protocol
[…]
[CAN bus data frame.]The picture shows an example of CAN bus data frame, with initial SOF, BaseID, IDE, Extended ID, RTR, Reserved, DLC, DATA from 0 to 64 bits, CRC, ACK, EOF
@@ 2.1.2. CAN vulnerabilities
Some approaches to secure the communication on the bus have been experimented during the years (Happel, 2014), but unfortunately they have not gained enough traction. Consequently, the protocol does not respect 2 of the 3 fundamental properties focused on by information security principles, which are summarized by the CIA triad, Confidentiality, Integrity, Availability:
@@
The most difficult task for an attacker, considering the critical vulnerabilities of the CAN protocol, would be to access the bus. The main attack vectors are two (Koscher et al., 2010), one is physical access, the other one is through one of the many wireless interfaces.
@@
A challenge-response mechanism protects ECUs from being reflashed and reading sensible memory. This challenge-response is performed through a seed-to-key challenge, but both seed and key are fixed and stored in the device. The system allows only an attempt every 10 seconds, but due to the low complexity, it is sufficient to have access for 7 and a half days to be able to crack the challenge-response and therefore reflash every ECU (Koscher et al., 2010).
The CAN bus exposes also a diagnostic system called Device Control, a powerful tool that is meant to be used for testing purposes, and it can modify the internal state of the ECUs. Paired with the vulnerabilities already described, it is easily exploitable and it can lead to dangerous consequences. This service takes an argument called a Control Packet Identifier (CPID), which specifies a group of controls to override, and each one can take up to five bytes as parameters, specifying which controls in the group are being overridden, and how to override them (Koscher et al., 2010).
@@ 2.2. Literature review and research gaps
One of the first security analyses on real cars has been made in 2010 by Koscher et al. (Koscher et al., 2010), assessing the behavior of the vehicle during specific attacks. They leveraged the security vulnerabilities of CAN to control a wide range of critical functions, like stopping the engine, disabling the brakes, or even braking on-demand individual wheels. However, they do not yet prove that these attacks are feasible remotely. In 2011 Checkoway et al. (Checkoway et al., 2011) analyzed purely the external attack surface of one vehicle to assess if it might be vulnerable to remote attacks exploiting the wireless interfaces present in the car, and they found out that it is indeed possible through multiple short-range and long-range vectors like Bluetooth or cellular interfaces. The results of Checkoway et al. are particularly concerning if we take into consideration the paper published by Koscher et al.: these two studies show that intruders can preserve their anonymity controlling critical features of vehicles from long distances, potentially causing disastrous, life-threatening consequences.
@@
Charlie Miller and Chris Valasek are two researchers working on automotive cybersecurity who gained a reputation during the years by publishing multiple relevant papers on the topic. In 2013 (Valasek and Miller, 2014) they tinkered with the CAN bus, exploiting its vulnerabilities to take control of critical functionalities of two different cars; in 2014 (Miller and Valasek, 2014) they took Checkoway et al.’s previous research further by identifying the remote attack surface for a large range of vehicles, identifying the ECUs’ capabilities and estimating the difficulty of remote exploitation; in 2015 (Miller and Valasek, 2015) they brought to conclusion 3 years of research disclosing a long-range remote attack that can be performed on Jeep Cherokees without any physical interaction, causing Fiat Chrysler to recall 1.4 million vehicles (News, 2015). One year later, in 2016 (Miller and Valasek, 2016), they published another paper explaining how they performed a masquerade attack on the Jeep Cherokee and how they were able to deal with message confliction.
@@
Charlie Miller and Chris Valasek believe that attack detection represents an inexpensive and accurate way to greatly improve the security of CAN networks (Miller and Valasek, 2014). Current research on the topic focuses extensively on machine learning (Bari et al., 2023) and deep learning (Lin et al., 2022) approaches in order to reach an increasingly higher accuracy. Unfortunately, these systems are not yet perfect: in 2021 Bhatia et al. (Bhatia et al., 2021) explained how they developed a masquerade attack called DUET which could evade detection of Voltage-based Intrusion Detection Systems. This type of IDS is able to observe the voltage fingerprint of ECUs, thus recognizing the ones belonging to attackers. However, the DUET attack showed that VIDS are not a silver bullet against intrusions, so the authors introduced a lightweight defense mechanism called RAID, suggesting to modify the frame format in a protocol-compatible way in order to safeguard against the corruption of ECUs’ voltage fingerprints.
@@
According to its design, an Intrusion Detection System can be categorized into two major families: anomaly-based and signature-based detection methods (Lokman et al., 2019). While signature-based IDSs compare the traffic with a database of known attack patterns, anomaly-based IDSs can be powered by machine learning techniques to model the normal behavior of the machine and they are able to detect unseen attacks as well.
@@
Although prior AI-based IDS models have yielded to promising results (Yang et al., 2022)(Hossain et al., 2020)(Song et al., 2020)(Nguyen et al., 2023), there exists a remarkable gap in testing them on a more realistic dataset including sophisticated attack patterns. It is worth noting that the effectiveness of an IDS model depends heavily on the quality of the data it was trained on. Previously introduced approaches were tested on datasets characterized by a limited number of attack types, typically only three or four, which are considered simple and unrealistic as they are conducted in a testbed environment (Lampe and Meng, 2023). These attacks are easily detectable and not necessarily verified on a real car. To build a robust IDS able to detect stealthy attacks it is imperative to test the model against high fidelity attack scenarios.
@@ 3.1. ROAD dataset and Labeling
The Real ORNL Automotive Dynamometer (ROAD) dataset (Verma et al., 2024) includes 33 attack captures spanning 30 minutes and 12 ambient captures over 3 hours (Table 1). It originates from a real undisclosed vehicle driven on a dynamometer of the Oak Ridge National Laboratory (ORNL) during 3 and a half hours of activity, reproducing various benign driving behaviors, while ambient and attack data were collected.
The logs are collected with the candump utility (CAN, 2015) and the output format is the following:
@@
    Targeted ID Fabrication Attacks: These attacks are performed injecting a message immediately after the legitimate message targeting a specific ID. In this case only the target signal is modified. Differently from most of public datasets, the fabrication attacks in the ROAD dataset contain one single frame injected between ambient frames of the same ID, since only a single injected message is needed after each legitimate message for this attack to be successful. Because these are real, physically verified attacks with the minimally occurring injected frames (due to the flam delivery), they provide perhaps the best (i.e., most stealthy/most difficult to detect), current, public data for testing frequency-based IDSs (Verma et al., 2024).
@@ 3.2. The HCRL Car-Hacking dataset
The Car-Hacking dataset (Hacking and Lab, 2018a), according to Rajapaksha et al. (2023) (Rajapaksha et al., 2023), is the most popular dataset in literature, and in many recent publications it seems to be used without citation (Verma et al., 2024). It is generated by injecting CAN messages into the Hyundai YF Sonata’s CAN bus through the OBD-II port and includes both attack and regular messages. Each sample is composed by a timestamp, the CAN ID, the DLC and up to 8 bytes of the data field. The attack types are four:
  • DoS attack: This attack is simply performed by transmitting messages on the CAN bus with the ID containing only dominant bits (0) (i.e. ID=0x000) every 0.3 milliseconds. As seen in the previous section, the CAN bus is using wired-AND signaling, which means that the messages with the ID containing the largest number of dominant bits will be prioritized over any other ECU, which will have to patiently wait for the attack frame to be completely transmitted before trying to send another one. Therefore, an attacker with access to the CAN bus would simply need to continuously inject messages using this technique and the correct functioning of the system would be disrupted, needing a manual restart of the car (Koscher et al., 2010).
@@ 3.3. The In-Vehicle Network Intrusion Detection Challenge dataset
The In-Vehicle Network Intrusion Detection Challenge dataset (Hacking and Lab, 2019) was used during the ”Information Security R&D Data Challenge 2019” and includes traffic data collected from three vehicles in stationary state: a Huyndai Sonata, a Kia Soul, and a Chevrolet Spark. We have used the datasets collected during the preliminary round, since they include three attack categories (Flooding, Fuzzy, and Malfunction) compared to the two attack categories of the final round. The attack types are the same of the Car-Hacking dataset, but with different names. The dataset is labelled and, like the Car-Hacking dataset, each sample is represented by a timestamp, a CAN ID, the DLC and the data field. For all three types of vehicles, the proportion of ”Flooding” attacks is relatively high, about 20% of the total messages, compared to only about 8% for the other two types (Table 3). The application of the same attack types and setup on three different vehicles makes this dataset unique and enables us to test transfer learning approaches on it, in order to improve the generalizability of the model.
@@ 4.1. LCCDE and ML approaches
To compare neural network approaches to traditional machine learning ones, gradient boosting and decision tree based approaches have also been tested on the available datasets. Specifically we considered Random Forest, LightGBM and the LCCDE model. These methods, like other machine learning algorithms, are preferable because of their speed, efficiency and explainability. Furthermore, the Leader Class and Confidence Decision Ensemble (LCCDE) (Yang et al., 2022), an ensemble model based on gradient boosting algorithms, was able to reach state-of-the-art results. The ensemble algorithm is divided into two stages: training and prediction. In the training stage, the three base models are trained on the labeled data. For each class, the leader is selected based on the best F1-score of the models, obtained from cross-validation. In case of a tie in F1-score, the model with the highest speed is chosen.
@@ 4.2. LSTM-based model
LSTMs are a type of Recurrent Neural Networks (RNNs) that are frequently and effectively applied for sequence classification tasks, therefore we have selected the state-of-the-art LSTM model described by Hossain et al. (Hossain et al., 2020) to test this approach on the previously specified datasets, since it achieved outstanding results on the Survival Analysis Dataset (Hacking and Lab, 2018b). The authors have tested many hyperparameters for both binary and multiclass classification. The best performing model was composed by a single LSTM layer with 512 units using single time steps and batches of size 512 as input. Each input sample includes ID, DLC (if available) and the bytes of the data field. We have performed multiclass classification on every dataset and binary classification only on the ROAD dataset, to compare its results with the DCNN and the TAN models. However, the approach of analyzing single messages can be considered vulnerable to replay attacks and since the authors used single time steps, the utility of a LSTM for this purpose is questionable.
Figure 2. Frame creation process of the DCNNv2.
[…]
[Frame creation process of the DCNNv2]The picture shows the improved process of how the grids are created and labeled: if an attack frame’s ID is present in the sequence, the frame is labeled as an attack (1), otherwise it is labeled as normal (0). Horizontally, the bits of the ID are positioned in the grid. It also shows the step=1 to create the sequences.
@@ 4.3. DCNN
The IDS model proposed by Song et al. (Song et al., 2020) is based on a convolutional neural network (CNN), aimed at detecting the patterns of IDs appearing in the CAN traffic. During the training phase, the model employs a frame builder to extract sequences of IDs from the logged traffic data. To facilitate CNN processing, which typically works with grid-formatted input, the temporal sequential patterns of CAN traffic are processed as spatially-local correlation. The frame builder extracts the 29-bit identifiers, which are the maximum number of bits for representing an ID when using the ’extended format’ (Semiconductor, 2017), from the most recent 29 messages and builds a $29\times 29$ bitwise frame by stacking them. A sliding window of size 29, with a step of 29 is used. This means that attack patterns might extend across 2 different frames and could not be easily recognized by the DCNN, which is unable to retain memory across frames. For this reason, we tested a different approach modifying the step of the sliding window to one, enabling the model to catch every pattern involving IDs and extending for any 29 consecutive frames. The frame creation process is illustrated in Figure 2.
@@
The DCNN model itself is based on the Inception-ResNet model (Szegedy et al., 2016), which is originally designed to classify $299\times 299\times 3$ input image data into 1000 classes, but has been reduced by the authors to classify more efficiently $29\times 29\times 1$ into only 2 classes. Compared to the original model the proposed DCNN excludes the Inception-ResNet-C block, while keeping the stem, Inception-ResNet-A, -B, Reduction-A and -B, average pooling, dropout, and softmax blocks, although they have been internally simplified to adapt to the studied case. The final model uses only 2% of the memory of the original one, and only the 18% of the original parameters, making it more suitable for use in restricted environments, like automotive applications.